The balance sheet for Sinking Ship Corp. is shown here in market value terms. There are 5,000 shares of stock outstanding.

  

Market Value Balance Sheet
  Cash $ 45,100   Equity $ 495,100
  Fixed assets 450,000
     Total $ 495,100      Total $ 495,100

  

Instead of a dividend of $1.40 per share, the company has announced a share repurchase of $7,000 worth of stock.

a.

How many shares will be outstanding after the repurchase? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. What will the price per share be after the repurchase? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a)current share price = equity /number of shares outstanding

                 = 495100/5000

                   = $ 99.02 per share

Number of shares repurchased = Amount paid to repurchase /current price

                                = 7000 / 99.02

                                = 70.69 shares

Shares after repurchase = current shares- repurchased shares

                                            = 5000-70.69

                                             = 4929.31 shares

b)shares price per share after repurchase = same as before = $ 99.02 per share

Calculations :

Equity after repurchase= 495100 - 7000 = 488100

price per share = 488100/4929.31

                = $ 99.02 per share